Grade 12Magical Mathematics[Interesting Approach]

In 7 power 1096 what is the digit in one`s place,in ten`s place,in hundreds place,plz give me complete explanation

lokesh palingi
12 Years agoGrade 12

1 Answer

Sher Mohammad
12 Years ago

(x+y)^n = {n \choose 0}x^n y^0 + {n \choose 1}x^{n-1}y^1 + {n \choose 2}x^{n-2}y^2 + \cdots + {n \choose n-1}x^1 y^{n-1} + {n \choose n}x^0 y^n,

3^0=1,3^1=3,3^2=9,3^3=27,3^4=81,

3^1096, will have 1 at one's place

3^1095, will have 7 at one's place'

3^1094, will have 9 at one's place

(10-3)^1096=3^1096-3^1095*10-......

at one's place it will have, 1

at 10' =0 at 100' 6.

Sher Mohammad,
B.Tech, IIT Delhi.